OpenStudy (ashy98):

Consider the function f(x) = 8 − 3x2. Find the value of f(x) when x = 3 and select the correct answer below. A.) −19 B.) −10 C.) 26 D.) 35

jimthompson5910 (jim_thompson5910):

Replace x with 3, then evaluate to get \[\Large f(x) = 8 - 3x^2\] \[\Large f(x) = 8 - 3(x)^2\] \[\Large f({\color{red}{x}}) = 8 - 3({\color{red}{x}})^2\] \[\Large f({\color{red}{3}}) = 8 - 3({\color{red}{3}})^2\] \[\Large f(3) = ???\]

OpenStudy (ashy98):

-19 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (ashy98):

yay thank you!!!(:

jimthompson5910 (jim_thompson5910):

you're welcome